40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 3997     Entry time: Tue Nov 30 12:45:36 2010
Author: kiwamu 
Type: Update 
Category: SUS 
Subject: found a loose connection : ITMX damping 

 Last night I found that the response of ITMX against the angle offsets were strage.

Eventually I found a loose connection at the feedthrough connectors of ITMX chamber.

 

So I pushed the connector hard, and then ITMX successfully became normal.

It looked like someone had accidentally kicked the cable during some works.

This bad connection had made unacceptable offsets in the OSEM readout, but now they seem fine.

ELOG V3.1.3-